Problem:
x^2+y^2=9
How to find answer y''= -x^2+y^2/y^3=-9/y^3 ?( I can't describe the operation type therefore any solution is accepted)
1
Expert's answer
2020-11-10T19:58:37-0500

Consider the equation:


"x^2+y^2=9" (*)


Find derivates of both parts of the equation (*):


"2x+2yy'=0" (**)


Further, let us find derivates of both parts of the equation (**):


"2+2(y')^2+2yy''=0" (***)


The equation (**) is equivalent to


"y'=-\frac{x}{y}" (****)


The equation (***) is equvialent to


"y''=-\frac{1+(y')^2}{y}="


 | next we use (****) |  


"=-\frac{1+(-\frac{x}{y})^2}{y}=-\frac{y^2+x^2}{y^3}="


 | next we use (*) |  


"=-\frac{9}{y^3}"